OpenStudy (anonymous):

Im pretty sure this might be simple but i tend to struggle with equations that include square roots so if anyone can help me thatd be lovely please...step by step if possible and explain if possible... Use Implicit Differentiation to find the derivative \[\sqrt{x}+\sqrt{y}=1\] please help...

zepdrix (zepdrix):

So for square root, recall that you can rewrite it with a rational exponent,\[\large\rm \sqrt{x}=x^{1/2}\]Applying power rule shows us how to differentiate this,\[\large\rm \frac{d}{dx}x^{1/2}\quad=\frac{1}{2}x^{-1/2}\quad=\frac{1}{2x^{1/2}}\quad=\frac{1}{2\sqrt x}\]

zepdrix (zepdrix):

But, I strongly recommend that you take the time to memorize this derivative. Square root shows up sooo often, just put it to memory. Derivative of square root is `one over two square roots` \[\large\rm \frac{d}{dx}\sqrt x=\frac{1}{2\sqrt x}\]

zepdrix (zepdrix):

\[\large\rm \sqrt{x}+\sqrt{y}=1\]So we're going to differentiate each side with respect to x.

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}\left(\sqrt{x}+\sqrt{y}\right)=\frac{d}{dx}1\]What do you get on the right side? :)

OpenStudy (anonymous):

i actually knew the rational exponent but never figured that the derivative went to \[\frac{ 1 }{ 2\sqrt{x} }\] thank you..and do you think that maybe you could possibly explain by differentiating y instead? its only because thats how i need to do it and i dont wish to become confused but if not its okay @zepdrix

zepdrix (zepdrix):

The `chain rule` is the most difficult of the basic differentiation rules to get a grasp of. Make sure you spend a lot of time practicing it. Realize, that will any problem, we can always apply the chain rule up to the variable of differentiation. Example:\[\large\rm \frac{d}{dx}(x)^2\]The power rule tells me how to differentiate, but I still need to apply chain rule to the inner function after that,\[\large\rm \frac{d}{dx}(x)^2\quad=2(x)^1\frac{d}{dx}x\]We can also think of this result in this way,\[\large\rm \frac{d}{dx}(x)^2\quad=2(x)^1\frac{dx}{dx}\]But the problem is, dx/dx has no real significance. It's measuring how much x changes, as x changes.\[\large\rm \frac{d}{dx}(x)^2\quad=2(x)^1(1)\]So we get this result, ya?\[\large\rm \frac{d}{dx}x^2=2x\]Hopefully this weird explanation isn't too long-winded :) But I feel this might help understand why we end up with these weird dy/dx looking things when we do implicit differentiation. So notice that when we differentiate something involving x, with respect to x, that very last chain is completely unnecessary.

zepdrix (zepdrix):

That is not the case when differentiating something involving y though.

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}\sqrt{x}\quad=\frac{1}{2\sqrt x}\frac{d}{dx}x\]\[\large\rm \frac{d}{dx}\sqrt{x}\quad=\frac{1}{2\sqrt x}\cdot 1\]

zepdrix (zepdrix):

Same process, but now this chain produces something important that doesn't simply go away.\[\large\rm \frac{d}{dx}\sqrt{y}=\frac{1}{2\sqrt{y}}\frac{d}{dx}y\]

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}\sqrt{y}=\frac{1}{2\sqrt{y}}\frac{dy}{dx}\]

zepdrix (zepdrix):

Questions? :) It can be a lot to take in heh

OpenStudy (anonymous):

No no you're good i am following what you are saying c: its after differentiating to the equation you displayed above that i have a problem isolating \[\frac{ dy }{ dx }\] to solve for it

zepdrix (zepdrix):

Oh i see :3

zepdrix (zepdrix):

Did you remember to differentiate the right side? Derivative of our constant 1 gives us 0, ya?

OpenStudy (anonymous):

Yes c:

zepdrix (zepdrix):

\[\large\rm \sqrt{x}+\sqrt{y}=1\]\[\large\rm \frac{1}{2\sqrt x}+\frac{1}{2\sqrt y}\frac{dy}{dx}=0\]So this is just basic Algebra from this point. We have a few different routes we can take.

zepdrix (zepdrix):

I don't like fractions, I would probably multiply both sides by the stuff in the denominators,\[\large\rm 2\sqrt x \sqrt y\left(\frac{1}{2\sqrt x}+\frac{1}{2\sqrt y}\frac{dy}{dx}\right)=0\]\[\large\rm \sqrt y+ \sqrt x \frac{dy}{dx}=0\]Then how be subtract sqrt y to the other side,\[\large\rm \sqrt x\frac{dy}{dx}=-\sqrt y\]And then divide the sqrt x over, ya? :)

zepdrix (zepdrix):

Then how bout* we* subtract sqrt y wow those typos today ;c

OpenStudy (anonymous):

So then the answer becomes \[\frac{ -\sqrt{y} }{ \sqrt{x}}\]?

zepdrix (zepdrix):

Yes, good. And if you like, you can combine the roots,\[\large\rm \frac{dy}{dx}=-\sqrt{\frac{y}{x}}\]

OpenStudy (anonymous):

Holy cheese its! thats amazing, you have no idea how thankful i am for your guidance because i honestly struggle so much in calculus because of this and tomorrow we have a test over implicit differentiation and im pretty sure square root will show up and i will be eternally grateful because i can understand the concept of it ^.^

zepdrix (zepdrix):

np, yay we did it \c:/

zepdrix (zepdrix):

tomorrow?? :O Oh boy, keep studying!

zepdrix (zepdrix):

Are you comfortable with your chain rule? Like, do you understand how to find the derivative of something like this?\[\large\rm \frac{d}{dx}\sqrt{x^2+2x}\]

OpenStudy (anonymous):

i know ive been studying the worksheets we received and it has the answers on it but i like to understand how a person gets the answer so i learn as well and well now i understand one of the problems c: Thanks to you

OpenStudy (anonymous):

And sadly...no. i failed that test horrendously so im still trying to work on it its only cause the class is fast paced and well she doesnt necessarily teach during tutoring either so i tend to ask here and there but still struggle for the most part with chain rule

zepdrix (zepdrix):

So we decided that the derivative of sqrt(x) is this,\[\large\rm \frac{d}{dx}\sqrt{x}\quad=\frac{1}{2\sqrt x}\]When the "stuff" under the root is more than just x, we `must` apply chain rule.\[\large\rm \frac{d}{dx}\sqrt{stuff}\quad=\frac{1}{2\sqrt{stuff}}\cdot(stuff)'\]multiplying by the derivative of the inner function.

OpenStudy (anonymous):

Okay im following slightly...proceed.

zepdrix (zepdrix):

So we apply square root derivative, and chain,\[\large\rm \frac{d}{dx}\sqrt{x^2+2x}=\frac{1}{2\sqrt{x^2+2x}}\cdot (x^2+2x)'\]The square root of the stuff turns into one over two square roots of the stuff. Chain rule tells us to make a copy of the inner function, take its derivative, and multiply. That prime that I wrote shows that we still need to take that derivative. The prime is sort of like ... "setting up" the chain. This should feel familiar if you've gone over Product and Quotient Rules at this point. "setting up" the derivative can be very helpful before actually taking them. To resolve the chain, we just power rule a couple times, ya?\[\large\rm \frac{d}{dx}\sqrt{x^2+2x}=\frac{1}{2\sqrt{x^2+2x}}\cdot (2x+2)\]

OpenStudy (anonymous):

Okay that makes sense c:

OpenStudy (anonymous):

Wait..so if were using power rule to find the derivative of the copy of the inner function do we also do the same to the original inner function(underneath the square root)?

zepdrix (zepdrix):

Nooo :O When differentiating "the outside", the inside of THAT doesn't change. Good question though.

OpenStudy (anonymous):

Okay cool, just verifying c:

OpenStudy (anonymous):

So...once you take the derivative of the "outside" what follows in order to actually solve the equation?

OpenStudy (anonymous):

Its the solving part that gets me because i dont necessarily like square root and fractions applied together...they're my weakness with implicit -.-

zepdrix (zepdrix):

Well this chain rule is producing these derivative things, y' or dy/dx if you like. With implicit differentiation, your goal is to solve for this y'. It's going to seem extra confusing when you have MULTIPLY y' floating around. Example:

zepdrix (zepdrix):

\[\large\rm \sqrt{x^2+2x}+y^2=y+1\]Differentiating with respect to x,\[\large\rm \frac{2x+2}{2\sqrt{x^2+2x}}+2yy'=y'+0\]The first term is just the example from before. Now you're expected to solve for this y'. You need to apply some fancy Algebra skills. Let's subtract 2yy' from each side,\[\large\rm \frac{2x+2}{2\sqrt{x^2+2x}}=y'-2yy'\]And then factor a y' out of each term on the right,\[\large\rm \frac{2x+2}{2\sqrt{x^2+2x}}=y'(1-2y)\]And then divide both sides by (1-2y),\[\large\rm \frac{2x+2}{2(1-2y)\sqrt{x^2+2x}}=y'\]That would be an example of "solving".

OpenStudy (anonymous):

Wow! o.O Thank you, how do you understand all of this? isnt it confusing because i know for me it truly is

zepdrix (zepdrix):

Haha :) I'm math major, I really truly miss this fun easy math. It becomes a lot less fun as you go on further XDD

zepdrix (zepdrix):

If you're not a math major, then it will probably always be a little bit weird. But after you've done tons and tons and tons of problems, these steps seem really simple after a while ^^

OpenStudy (anonymous):

Youre a Math major? Man no wonder this is so simple to you XD I still got a long way for the harder stuff to come but yes! Alot of practice it shall take

zepdrix (zepdrix):

I have another example just in case you'd like to see :3

zepdrix (zepdrix):

Example 2:\[\large\rm \sqrt{x^2+y^2}=y-1\]Differentiating, setting up our chain rule,\[\large\rm \frac{1}{2\sqrt{x^2+y^2}}(x^2+y^2)'=y'-0\]resolving chain rule gives us,\[\large\rm \frac{1}{2\sqrt{x^2+y^2}}(2x+2yy')=y'-0\]Divide the 2's out maybe,\[\large\rm \frac{1}{\sqrt{x^2+y^2}}(x+yy')=y'-0\]Distribute the fraction to each term in the brackets,\[\large\rm \frac{x}{\sqrt{x^2+y^2}}+\frac{y}{\sqrt{x^2+y^2}}y'=y'\]Again let's subtract the (stuff)y' term to the other side,\[\large\rm \frac{x}{\sqrt{x^2+y^2}}=y'-\frac{y}{\sqrt{x^2+y^2}}y'\]Again, factor out the y',\[\large\rm \frac{x}{\sqrt{x^2+y^2}}=y'\left(1-\frac{y}{\sqrt{x^2+y^2}}\right)\]And then divide by that big mess to isolate your y',\[\large\rm \frac{x}{\left(1-\frac{y}{\sqrt{x^2+y^2}}\right)\sqrt{x^2+y^2}}=y'\]And obviously that answer can be written in a much much nicer way, but let's not worry about that.

zepdrix (zepdrix):

Examples are good c: heh

OpenStudy (anonymous):

Examples are rich i must say c: An for you to actually give so many makes you extremely pro

OpenStudy (anonymous):

But i must say that you seem to enjoy this c: Thank god you helped me

zepdrix (zepdrix):

"No crying in calculus" lol silly teacher :3

OpenStudy (anonymous):

Yes...she is very fond of that quote because everytime she teaches us something that seems easy even though we know its not we always tell her "thats it im going to cry.." and well yeah she laughs because she's been there but because shes also a math major well...she finds it silly that we complain XD

zepdrix (zepdrix):

hah

OpenStudy (anonymous):

Well i would like to thank you again for the help and examples you have given me but i shall now try and get some rest for the *uses sarcasm* Amazing test that i am so looking forward to tomorrow or later technically...but eh Anyways, thank you and hopefully if i happen to have another question about something i can go to you for help again c: But yeah

zepdrix (zepdrix):

ok np \c: have good night

OpenStudy (anonymous):

I shall, Good Night and thanks a million times again hah